# SemGauss-SLAM: Dense Semantic Gaussian Splatting SLAM

### v2

@@ Abstract
We propose SemGauss-SLAM, the first semantic SLAM system utilizing 3D Gaussian representation, that enables accurate 3D semantic mapping, robust camera tracking, and high-quality rendering in real-time. In this system, we incorporate semantic feature embedding into 3D Gaussian representation, which effectively encodes semantic information within the spatial layout of the environment for precise semantic scene representation. Furthermore, we propose feature-level loss for updating 3D Gaussian representation, enabling higher-level guidance for 3D Gaussian optimization. In addition, to reduce cumulative drift and improve reconstruction accuracy, we introduce semantic-informed bundle adjustment leveraging semantic associations for joint optimization of 3D Gaussian representation and camera poses, leading to more robust tracking and consistent mapping. Our SemGauss-SLAM method demonstrates superior performance over existing dense semantic SLAM methods in terms of mapping and tracking accuracy on Replica and ScanNet datasets, while also showing excellent capabilities in novel-view semantic synthesis and 3D semantic mapping. The source code will be released upon acceptance.
Keywords:
Dense Semantic SLAM Semantic-informed Bundle Adjustment 3D Gaussian Semantic Representation 3D Reconstruction
@@ 1 Introduction
Dense semantic Simultaneous Localization and Mapping (SLAM) is a fundamental challenge for robotic systems [18, 15] and autonomous driving [1, 10]. It integrates semantic understanding of the environment into dense map reconstruction and performs pose estimation simultaneously. Traditional semantic SLAM has limitations including its inability to predict unknown areas and the demand for considerable map storage [11]. Subsequent semantic SLAM based on Neural Radiance Fields (NeRF) [16] methods [32, 9] have addressed these drawbacks, but suffers from inefficient per-pixel raycasting rendering.
Recently, a novel radiance field based on 3D Gaussian [8] has demonstrated remarkable capability in scene representation, enabling high-quality and efficient rendering via splatting. Following the advantages of 3D Gaussian representation, 3D Gaussian SLAM methods [27, 7, 31, 13] have been developed to achieve photo-realistic mapping. However, existing 3D Gaussian SLAM systems focus on visual mapping to obtain RGB maps, where color information alone is insufficient for downstream tasks such as navigation. Meanwhile, current semantic SLAM methods based on NeRF are prone to cumulative drift, leading to degraded SLAM accuracy. Therefore, developing a low-drift semantic SLAM system based on radiance field is essential and challenging.
@@
For radiance field-based semantic SLAM, there are two challenges: 1) Existing semantic SLAM methods struggle to achieve real-time 3D semantic mapping from 2D semantic information without predefined bounds, which is a fundamental requirement for semantic SLAM system. 2) As the progression of tracking, pose estimation is prone to cumulative drift, leading to a decrease in SLAM accuracy.
For the first challenge, SNI-SLAM [32] utilizes feature collaboration to achieve semantic mapping but requires specific scene bounds. Moreover, solely utilizing rendering results for semantic representation optimization without initial value is computationally inefficient. In this paper, we leverage the explicit structure of 3D Gaussian representation and propagate extracted 2D semantic features directly to 3D Gaussian as the initial value for achieving unbounded mapping and efficient 3D semantic scene optimization. This design enables faster transfer from 2D semantic information to 3D semantic map. In addition, we introduce 2D feature-level loss to guide scene representation optimization at a higher-level, thereby accelerating the convergence of 3D Gaussian optimization.
For the second challenge, NICE-SLAM [33] performs local bundle adjustment (BA) for joint optimization of pose and scene representation to reduce accumulated drift. However, it only utilizes appearance and geometry constraints, lacking adequate semantic constraints for semantic SLAM optimization. To address this challenge, we leverage semantic associations among co-visible frames and propose semantic-informed bundle adjustment for joint optimization of camera poses and 3D Gaussian representation. This design exploits the consistency of multi-view semantics for establishing constraints, which enables the reduction of cumulative drift in tracking and enhanced semantic mapping precision.
Overall, we provide the following contributions:
  • We present SemGauss-SLAM, the first 3D Gaussian semantic SLAM system, which can achieve accurate semantic mapping and photo-realistic reconstruction. We incorporate semantic feature embedding into 3D Gaussian for precisely constructing semantic maps. Moreover, feature-level loss is introduced to provide higher-level guidance for 3D Gaussian optimization.
@@
  • We conduct extensive evaluations on two challenging datasets, Replica [21] and ScanNet [2], to demonstrate our method achieves state-of-the-art performance compared with existing radiance field-based SLAM in mapping, tracking, semantic segmentation and novel-view synthesis.
@@ 2 Related Work
Semantic SLAM can be divided into two main categories based on the form of scene representation: traditional semantic SLAM and neural implicit semantic SLAM. Traditional semantic SLAM employs explicit 3D representation such as surfels [15], mesh [18, 24, 5] and Truncated Signed Distance Fields (TSDF) [20, 14, 3]. Neural implicit semantic SLAM [4, 9, 32] utilizes implicit representation, such as feature grid [9] and feature plane [32].
@@
Traditional Semantic SLAM.   SemanticFusion [15] uses surfel representation and employs Conditional Random Field (CRF) for updating class probability distribution incrementally. Fusion++ [14] performs object-level SLAM where each object is reconstructed within its own TSDF volume and segmented based on estimated foreground probability. Kimera [18] utilizes visual-inertial odometry for pose estimation and generates dense semantic mesh maps. However, these explicit scene modeling methods not only require high storage space but also fail to achieve high-fidelity and complete reconstruction.
@@
Neural Implicit SLAM.   iMAP [23] first achieves real-time mapping and tracking utilizing a single MLP network for scene representation. To overcome oversmoothed scene reconstruction and improve scalability, NICE-SLAM [33] adopts hierarchical feature grid representation. Following this, several works [29, 6, 25, 19] introduce more efficient scene representation, such as hash-based feature grid and feature plane, to achieve more accurate SLAM performance.
@@
For neural implicit semantic SLAM, DNS SLAM [9] utilizes 2D semantic priors and integrates multi-view geometry constraints for semantic reconstruction. SNI-SLAM [32] introduces feature collaboration and one-way correlation decoder for improved scene representation in semantic mapping. However, these methods require specific scene bounds for mapping and suffer from cumulative drift in pose estimation. In this paper, we leverage the explicit structure of 3D Gaussian for unbounded mapping, and perform semantic-informed bundle adjustment utilizing multi-frame semantic constraints for conducting low-drift, high-quality dense semantic SLAM.
@@
3D Gaussian Splatting SLAM.   3D Gaussian [8] emerges as a promising 3D scene representation using a set of 3D Gaussians with position, anisotropic covariance, opacity, and color. This representation is capable of quick differential rendering through splatting and has a wide range of applications in dynamic scene modeling [12, 26, 28] and scene editing [34, 30].
@@
Our main focus is on 3D Gaussian SLAM [27, 7, 13, 31]. These works emerge concurrently and all perform dense visual SLAM by leveraging scene geometry and appearance modeling capabilities of 3D Gaussian representation. SplaTAM [7] introduces silhouette-guided optimization to facilitate structured map expansion for dense mapping of visual SLAM. Gaussian Splatting SLAM [13] performs novel Gaussian insertion and pruning for monocular SLAM. However, the capability of 3D Gaussian representation extends well beyond appearance and geometry modeling of scene, as it can be augmented to perform semantic scene understanding. In this paper, we incorporate semantic feature embedding into 3D Gaussian for 3D semantic scene modeling to achieve high-precision dense semantic SLAM.
@@
Figure 2: An overview of SemGauss-SLAM. Our method takes an RGB-D stream as input. RGB images are fed into feature extractor to obtain semantic features. These features are then categorized by pretrained classifier to attain semantic labels. Then, semantic features, semantic labels, along with the input RGB and depth data serve as supervision signals. In the meantime, semantic features and input RGB-D data propagate to 3D Gaussian blobs as initial properties of Gaussian representation. Rendered semantic feature, RGB, and depth are obtained from 3D Gaussian splatting, while rendered semantic label is attained by classifying rendered feature. Supervision and rendered information are utilized for loss construction to optimize camera poses and 3D Gaussian representation. During the SLAM process, we utilize semantic-informed bundle adjustment based on multi-frame constraints for joint optimization of poses and 3D Gaussian representation.
@@ 3 Method
SemGauss-SLAM is the first RGB-D dense semantic SLAM based on 3D Gaussian splatting. The overview of SemGauss-SLAM is shown in Fig. 2. Sec. 3.1 introduces semantic Gaussian representation for dense semantic mapping, as well as the process of 3D Gaussian-based semantic SLAM. Sec. 3.2 introduces the loss functions in the mapping and tracking process. Sec. 3.3 presents semantic-informed bundle adjustment to reduce accumulated drift in tracking and improve semantic mapping accuracy.
(a) 3D Gaussian Semantic Mapping
(b) Semantic-informed Bundle Adjustment
Figure 3: (a) 3D Gaussian semantic mapping process that utilizes 3D Gaussian blobs with semantic feature embedding for visualization. (b) Illustration of semantic-informed bundle adjustment.
@@
Semantic Gaussian Representation.   We utilize a set of Gaussians with specific properties for scene representation. To simplify scene representation for the SLAM process, we use isotropy Gaussians. Considering semantic representation, which has never been explored in other Gaussian SLAM, we introduce a new parameter, semantic feature embedding, to each Gaussian for semantic representation. To achieve real-time semantic mapping, it is crucial for 3D Gaussian representation, augmented by semantic feature embedding, to converge rapidly during the mapping optimization process. Therefore, instead of initializing these feature embeddings randomly, we propagate 2D semantic features extracted from images to 3D Gaussian as the initial values to achieve faster convergence of Gaussian semantic feature optimization. Moreover, integrating feature embedding into 3D Gaussian makes Gaussian semantic representation compact and efficient for capturing the spatial semantic information of the environment. Overall, each Gaussian includes 3D center position $\mu$, radius $r$, color $c=(r,g,b)$, opacity $\alpha$, 16-channel semantic feature embedding $e$, and is defined as standard Gaussian equation multiplied by opacity $\alpha$:
@@
3D Gaussian Rendering.   Following Gaussian splatting [8], we project 3D Gaussians to 2D splats for high-fidelity differentiable rendering, allowing for explicit gradient flow for Gaussian scene optimization and pose estimation. Specifically, for splatting of semantic feature embedding, we first sort all Gaussians from front-to-back and then blend $N$ ordered points projecting to pixel $p=(u,v)$, obtaining 2D semantic feature maps $E(p)$:
@@
where $C(p)$ and $D(p)$ represents splatted 2D color and depth images respectively. Moreover, given camera pose, visibility information of Gaussian is required for mapping and tracking process, such as adding new Gaussian and loss construction. Therefore, following  [7], we render a silhouette image to determine visibility:
@@
Tracking Process.   During tracking, we keep 3D Gaussian parameters fixed and only optimize camera pose $T$ of current frame. Since the proximity between adjacent frames is relatively small, we assume a constant velocity model to obtain an initial pose estimation for new frame. The camera pose is then iteratively refined by optimization through loss construction between differentiable rendering from 3D Gaussian and camera observation within the visible silhouette.
Mapping Process.   Our system performs RGB mapping and semantic mapping simultaneously. Mapping process begins with the initialization of the scene representation, which is achieved by inverse transforming all pixels of the first frame to 3D coordinates and obtaining the initial 3D Gaussian representation. Then, when the overlap of coming frame with existing map rendering is less than half, we add a new Gaussian for incremental mapping. Figure 3(a) demonstrates 3D Gaussian semantic mapping process.
@@ 3.2 Loss Functions
For optimization of semantic scene representation, we utilize cross-entropy loss for constructing semantic loss $\mathcal{L}_{s}$. Furthermore, instead of using semantic loss solely, we introduce a feature-level loss $\mathcal{L}_{f}$ for higher-level semantic optimization guidance. Feature-level loss is obtained by constructing L1 loss between extracted features generated by Dinov2 [17]-based feature extractor and splatted features obtained from 3D Gaussian representation. Compared with semantic loss, feature loss provides direct guidance on intermediate features, leading to more robust and accurate semantic scene understanding.
@@
In mapping process, we construct loss over all rendered pixels for 3D Gaussian scene optimization. Moreover, we add SSIM term to RGB loss following [8]. The complete loss function for mapping is a weighted sum of the above losses:
$\mathcal{L}_{\text{mapping}}=\sum_{p\in P_{M}}{(\lambda_{f_{m}}\mathcal{L}_{f}(p)+\lambda_{s_{m}}\mathcal{L}_{s}(p)+\lambda_{c_{m}}\mathcal{L}_{c}(p)+\lambda_{d_{m}}\mathcal{L}_{d}(p))},$ (6)
where $P_{M}$ represents the collection of all pixels in rendered image. $\lambda_{f_{m}}$, $\lambda_{s_{m}}$, $\lambda_{c_{m}}$, $\lambda_{d_{m}}$ are weighting coefficients.
During tracking, utilizing an overly constrained loss function for pose estimation can lead to less precise camera pose and increased processing time, which is unsuitable for real-time tracking. Therefore, loss function for tracking is constructed based only on a weighted sum of RGB loss and depth loss:

### v3

@@ Abstract
We propose SemGauss-SLAM, a dense semantic SLAM system utilizing 3D Gaussian representation, that enables accurate 3D semantic mapping, robust camera tracking, and high-quality rendering simultaneously. In this system, we incorporate semantic feature embedding into 3D Gaussian representation, which effectively encodes semantic information within the spatial layout of the environment for precise semantic scene representation. Furthermore, we propose feature-level loss for updating 3D Gaussian representation, enabling higher-level guidance for 3D Gaussian optimization. In addition, to reduce cumulative drift in tracking and improve semantic reconstruction accuracy, we introduce semantic-informed bundle adjustment leveraging multi-frame semantic associations for joint optimization of 3D Gaussian representation and camera poses, leading to low-drift tracking and accurate mapping. Our SemGauss-SLAM method demonstrates superior performance over existing radiance field-based SLAM methods in terms of mapping and tracking accuracy on Replica and ScanNet datasets, while also showing excellent capabilities in high-precision semantic segmentation and dense semantic mapping.
@@ 1 Introduction
Dense semantic Simultaneous Localization and Mapping (SLAM) is a fundamental challenge for robotic systems rosinol2020kimera ; mccormac2017semanticfusion and autonomous driving bao2022semantic ; lianos2018vso ; wang2022sfgan ; jiang20223d ; jiang20243dsflabelling ; jiang2022ffpa ; feng2024rogs . It integrates semantic understanding of the environment into dense map reconstruction and performs pose estimation simultaneously. Traditional dense semantic SLAM has limitations including its inability to predict unknown areas liu2023unsupervised . Subsequent semantic SLAM based on Neural Radiance Fields (NeRF) mildenhall2021nerf methods zhu2023sni ; li2023dns ; haghighi2023neural have addressed these drawbacks, but suffer from inefficient per-pixel raycasting rendering for real-time optimization of implicit scene representation and low-quality novel view semantic representation.
Recently, a novel radiance field based on 3D Gaussian Splatting (3DGS) kerbl20233d has demonstrated remarkable capability in scene representation, enabling high-quality and efficient rendering via splatting. Following the advantages of 3D Gaussian representation, 3DGS-based SLAM methods yan2023gs ; keetha2023splatam ; yugay2023gaussian ; matsuki2023gaussian ; huang2023photo have been developed to achieve photo-realistic mapping. However, most existing 3DGS SLAM systems focus on visual mapping to obtain RGB maps, where color information alone is insufficient for downstream tasks such as navigation. Moreover, current semantic NeRF-based SLAM methods suffer from cumulative drift in tracking, leading to degraded SLAM accuracy. Additionally, these methods struggle to achieve accurate semantic segmentation from novel viewpoints, indicating limited precision in online-reconstructed 3D semantic representation. Therefore, developing a low-drift and high-precision semantic SLAM system based on radiance field is essential and challenging.
@@
To overcome this challenge, we introduce a novel dense semantic SLAM system based on 3D Gaussian Splatting with semantic-informed bundle adjustment, which enables high accuracy in tracking, dense mapping, and semantic segmentation from novel view. Different from concurrent work SGS-SLAM li2024sgs , which uses color for semantic representation to achieve 3DGS semantic SLAM, our method integrates semantic feature embedding into 3D Gaussian for semantic modeling, resulting in a more discriminative and comprehensive understanding of the scene. Moreover, by leveraging the explicit 3D structure of 3D Gaussian representation, our embedding-based semantic Gaussian representation is capable of capturing the spatial distribution of semantic information, achieving high-accuracy novel view semantic segmentation. Furthermore, to reduce accumulated drift in radiance field-based semantic SLAM, we leverage semantic associations among co-visible frames and propose semantic-informed bundle adjustment (BA) for joint optimization of camera poses and 3D Gaussian representation. This design exploits the consistency of multi-view semantics for establishing constraints, which enables the reduction of cumulative drift in tracking and enhanced semantic mapping precision. Overall, we provide the following contributions:
  • We present SemGauss-SLAM, a dense semantic SLAM system based on 3D Gaussian, which can achieve accurate semantic mapping, photo-realistic reconstruction, and robust tracking. We incorporate semantic feature embedding into 3D Gaussian to achieve precise semantic scene representation. Moreover, feature-level loss is introduced to provide higher-level guidance for 3D Gaussian optimization, leading to high-precision scene optimization results.
@@
  • We conduct extensive evaluations on two challenging datasets, Replica straub2019replica and ScanNet dai2017scannet , to demonstrate our method achieves state-of-the-art performance compared with existing radiance field-based SLAM in mapping, tracking, semantic segmentation and novel-view synthesis.
@@ 2 Related Work
Traditional Semantic SLAM.   Traditional semantic SLAM employs explicit 3D representation such as surfels mccormac2017semanticfusion , mesh rosinol2020kimera ; tian2022kimera ; hughes2022hydra and Truncated Signed Distance Fields (TSDF) schmid2022panoptic ; mccormac2018fusion++ ; grinvald2019volumetric for dense semantic mapping. SemanticFusion mccormac2017semanticfusion uses surfel representation and employs Conditional Random Field (CRF) for updating class probability distribution incrementally. Fusion++ mccormac2018fusion++ performs object-level SLAM where each object is reconstructed within its own TSDF volume and segmented based on estimated foreground probability. Kimera rosinol2020kimera utilizes visual-inertial odometry for pose estimation and generates dense semantic mesh maps. However, these explicit scene modeling methods not only require high storage space but also fail to achieve high-fidelity and complete 3D reconstruction.
@@
Neural Implicit SLAM.   Neural implicit representation mildenhall2021nerf has shown promising capability in scene reconstruction for dense mapping. iMAP imap first achieves real-time mapping and tracking utilizing a single MLP network for scene representation. To overcome over-smoothed scene reconstruction and improve scalability, NICE-SLAM zhu2022nice adopts hierarchical feature grid representation. Following this, several works yang2022vox ; eslam ; wang2023co ; sandstrom2023point ; kong2023vmap ; hu2024cp ; deng2023plgslam introduce more efficient scene representation, such as hash-based feature grid and feature plane, to achieve more accurate SLAM performance.
@@
For neural implicit semantic SLAM, DNS SLAM li2023dns utilizes 2D semantic priors and integrates multi-view geometry constraints for semantic reconstruction. SNI-SLAM zhu2023sni introduces feature collaboration and one-way correlation decoder for improved scene representation in semantic mapping. However, these methods require specific scene bounds for mapping and suffer from cumulative drift in pose estimation. In this paper, we leverage the explicit structure of 3D Gaussian for unbounded mapping and perform semantic-informed bundle adjustment utilizing multi-frame semantic constraints for conducting low-drift and high-quality dense semantic SLAM.
@@
3D Gaussian Splatting SLAM.   3D Gaussian Splatting kerbl20233d emerges as a promising 3D scene representation using a set of 3D Gaussians with learnable properties, including 3D center position, anisotropic covariance, opacity, and color. This representation is capable of quick differential rendering through splatting and has a wide range of applications in dynamic scene modeling luiten2023dynamic ; wu20234d ; yang2023deformable ; yan2024street ; zhou2023drivinggaussian , semantic segmentation cen2023segment ; ye2023gaussian ; zhou2023feature , scene editing zhuang2024tip ; fang2023gaussianeditor ; huang2023point .
@@
Our main focus is on 3D Gaussian SLAM yan2023gs ; keetha2023splatam ; matsuki2023gaussian ; yugay2023gaussian . These works emerge concurrently and all perform dense visual SLAM by leveraging scene geometry and appearance modeling capabilities of 3D Gaussian representation. SplaTAM keetha2023splatam introduces silhouette-guided optimization to facilitate structured map expansion for dense mapping of visual SLAM. Gaussian Splatting SLAM matsuki2023gaussian performs novel Gaussian insertion and pruning for monocular SLAM. However, the capability of 3D Gaussian representation extends well beyond appearance and geometry modeling of the scene, as it can be augmented to perform semantic scene understanding. The concurrent work SGS-SLAM li2024sgs introduces semantic color associated with the Gaussian for semantic representation. However, this color-based semantic modeling method ignores the higher-level information inherent in semantics, which is insufficient for semantic representation of the environment. To address this limitation, we incorporate semantic feature embedding into 3D Gaussian for 3D semantic scene modeling to achieve high-precision dense semantic SLAM.
@@
Figure 2: An overview of SemGauss-SLAM. Our method takes an RGB-D stream as input. RGB images are fed into feature extractor to obtain semantic features. These features are then categorized by a pretrained classifier to attain semantic labels. Then, semantic features, semantic labels, along with the input RGB and depth data serve as supervision signals. In the meantime, semantic features and input RGB-D data propagate to 3D Gaussian blobs as initial properties of Gaussian representation. Rendered semantic features, RGB, and depth are obtained from 3D Gaussian splatting, while rendered semantic labels are attained by classifying rendered features. Supervision and rendered information are utilized for loss construction to optimize camera poses and 3D Gaussian representation. During the SLAM process, we utilize semantic-informed bundle adjustment based on multi-frame constraints for joint optimization of poses and 3D Gaussian representation.
@@ 3 Method
The overview of SemGauss-SLAM is shown in Fig. 2. Sec. 3.1 introduces semantic Gaussian representation for dense semantic mapping, as well as the process of 3DGS-based semantic SLAM. Sec. 3.2 introduces loss functions in the mapping and tracking process. Sec. 3.3 presents semantic-informed bundle adjustment to reduce accumulated drift in tracking and improve mapping accuracy.
@@
Semantic Gaussian Representation.   We utilize a set of isotropy Gaussians with specific properties for scene representation in SLAM. To achieve semantic Gaussian mapping, we introduce a new parameter, semantic feature embedding, to each Gaussian for semantic representation. This design enables compact and efficient Gaussian semantic representation to capture the spatial semantic information of the environment. Moreover, it is crucial for 3D Gaussian representation, augmented by semantic feature embedding, to converge rapidly during optimization process for real-time mapping. Therefore, instead of initializing these feature embeddings randomly, we propagate 2D semantic features extracted from images to 3D Gaussian as the initial values to achieve faster convergence of semantic Gaussian optimization. In this work, we use a universal feature extractor Dinov2 oquab2023dinov2 , followed by a pretrained classifier to construct the segmentation network. Overall, each Gaussian includes 3D center position $\mu$, radius $r$, color $c=(r,g,b)$, opacity $\alpha$, 16-channel semantic feature embedding $e$, and is defined as standard Gaussian equation multiplied by opacity $\alpha$:
@@
3D Gaussian Rendering.   Following Gaussian splatting kerbl20233d , we project 3D Gaussians to 2D splats for high-fidelity differentiable rendering, which facilitates explicit gradient flow for Gaussian scene optimization and pose estimation. Specifically, for splatting of semantic feature embedding, we first sort all Gaussians in order of depth from near to far and then blend $N$ ordered points projecting to pixel $p=(u,v)$, obtaining 2D semantic feature maps $E(p)$:
@@
where $C(p)$ and $D(p)$ represents splatted 2D color and depth images respectively. Moreover, given the camera pose, visibility information of Gaussian is required for mapping and tracking process, such as adding new Gaussian and loss construction. Therefore, following  keetha2023splatam , we render a silhouette image to determine visibility:
@@
Tracking Process.   During tracking, we keep 3D Gaussian parameters fixed and only optimize camera pose $T$ of current frame. Since the proximity between adjacent frames is relatively small, we assume a constant velocity model to obtain an initial pose estimation for the new frame. The camera pose is then iteratively refined by optimization through loss construction between differentiable rendering from 3D Gaussian and camera observation within the visible silhouette.
Mapping Process.   Our system performs RGB mapping and semantic mapping simultaneously. Mapping process begins with the initialization of the scene representation, which is achieved by inverse transforming all pixels of the first frame to 3D coordinates and obtaining the initial 3D Gaussian representation. Then, when the overlap of coming frame with rendering of the existing map is less than half, we add a new Gaussian for incremental mapping.
@@ 3.2 Loss Functions
For optimization of semantic scene representation, we utilize cross-entropy loss for constructing semantic loss $\mathcal{L}_{s}$. Furthermore, instead of using semantic loss solely, we introduce a feature-level loss $\mathcal{L}_{f}$ for higher-level semantic optimization guidance. Feature-level loss is obtained by constructing L1 loss between extracted features generated by Dinov2 oquab2023dinov2 -based feature extractor and splatted features obtained from 3D Gaussian representation. Compared with semantic loss, feature loss provides direct guidance on intermediate features, leading to more robust and accurate semantic scene understanding.
@@
In mapping process, we construct loss over all rendered pixels for 3D Gaussian scene optimization. Moreover, we add SSIM term to RGB loss following kerbl20233d . The complete loss function for mapping is a weighted sum of the above losses:
$\mathcal{L}_{\text{mapping}}=\sum_{p\in P_{M}}{(\lambda_{f_{m}}\mathcal{L}_{f}(p)+\lambda_{s_{m}}\mathcal{L}_{s}(p)+\lambda_{c_{m}}\mathcal{L}_{c}(p)+\lambda_{d_{m}}\mathcal{L}_{d}(p))}.$ (6)
$P_{M}$ represents a set of all pixels in rendered image. $\lambda_{f_{m}}$, $\lambda_{s_{m}}$, $\lambda_{c_{m}}$, $\lambda_{d_{m}}$ are weighting coefficients.
During tracking, utilizing an overly constrained loss function for pose estimation can lead to decreased accuracy in camera pose and increased processing time. Therefore, loss function for tracking is constructed based only on a weighted sum of RGB loss and depth loss: